# Source reconstruction algorithms for coupled parabolic systems from internal measurements of one scalar state

### v3

@@ Abstract.
This paper is devoted to the study of source reconstruction algorithms for coupled systems of heat equations, with either constant or spatially dependent coupling terms, where internal measurements are available from a reduced number of observed states. Two classes of systems are considered. The first comprises parabolic equations with constant zero-order coupling terms (through a matrix potential term or via the diffusion matrix). The second type considers parabolic equations coupled by a matrix potential that depends on spatial variables, which leads to the analysis of a non-self-adjoint operator. In all configurations, the source is assumed to be of separate variables, the temporal part is a known scalar function, and the spatial dependence is an unknown vector field. Several numerical examples using the finite element method in 1D and 2D are presented to show the reconstruction of space-dependent sources.
@@
Key words and phrases:
@@ 1.1. Introduction
  • The case where the coupling appears in the principal part:
    $\left\{\begin{array}[]{llll}\partial_{t}Y+DAY=\sigma(t)F(x),&\mbox{in}&\Omega\times(0,T),\\ Y(\,\cdot\,,0)=0,&\mbox{in}&\Omega,\end{array}\right$. (1.2)
    with $D\in\mathcal{M}_{n}(\mathbb{K})$ a diffusion matrix with constant coefficients. We assume that $D$ is diagonalizable with positive eigenvalues. In this setting, our question is: can we recover the source term $F=(f_{1},\dots,f_{n})^{t}$ in system (1.2) from a reduced number of measurements of the solution in $\mathcal{O}\times(0,T)$?
@@ 1.2. State of the art
    Inverse coefficient problems in coupled systems. Under the structure of systems (1.3) with $A=-\Delta$, the Bukhgeim-Klibanov method [32], which is based on Carleman estimates, has been widely employed to obtain results for inverse coefficient problems. For instance, the articles [16, 10] proved Lipschitz-type stability inequalities for $2\times 2$ reaction-diffusion systems using a single observation in a subdomain. These results were further extended to a nonlinear case in [17]. In a different context, the recent work [13] addressed the uniqueness and Lipschitz-type stability for cascade hyperbolic systems composed of $n$ equations, where internal measurements were available for all components of the solution except the last one. Another recent work [20] considered two coupled Schrödinger equations (with $A=i\Delta$ in (1.3)), in which the authors investigated the logarithmic stability for determining two potentials from internal observations of a single component of the solution. In contrast to the previous results, [20] combined Carleman estimates along with the Fourier-Bros-Iagolnitzer transform in order to prove the logarithmic stability. The paper [15] tackled the problem of identification for two discontinuous coefficients in a one-dimensional coupled parabolic system, observing only one component. The authors proved theoretical results based on Carleman-type inequalities, while numerical results were obtained using the finite difference method for both temporal and spatial discretization, jointly with the interior-point method (meaning that the numerical part is treated as an optimization problem). To summarize, these articles show identifiability and stability results for hyperbolic and parabolic systems coupled through potential matrices (i.e., systems as (1.3)).
    On the other hand, relatively few works have addressed inverse problems involving coupling through the principal term of the operator as in (1.2). The recent paper [14] addressed the coefficient determination problem for a transmitted disease model (coupled parabolic equations related to the SIR model) whose coupling is located in the principal term associated with the operator. The authors in [14] applied optimal control techniques with constraints to establish their main results. Additionally, the recent article [41], proved Hölder-type stability results (via Carleman inequalities) for the inverse coefficient problem based on internal observation data for a three-dimensional system of two coupled heat equations, similar to the structure given in (1.2).
@@
The remaining of the paper is organized as follows: Section 3 provides explicit source reconstruction formulas for coupled parabolic systems with constant coefficient matrices, either in the principal part of the operator or as a potential term, thus solving Problem 1. Section 4 addresses Problem 2, presenting a reconstruction formula for a one-dimensional $2\times 2$ parabolic system coupled through a non-constant zero-order matrix potential. Here, the main challenge arises from the non-self-adjoint operator, necessitating careful use of eigenfunctions associated with non-self-adjoint Sturm–Liouville operators to achieve reconstruction based solely on interior measurements and a reduced number of states. Section 5 validates these theoretical findings through numerical experiments implemented with finite element methods and gradient descent algorithms. These experiments include reconstructions for Problem 2 in 1D, analyses of sensitivity to regularization parameters and observational noise, and reconstructions addressing Problem 1 in 2D. Finally, the Lipschitz-type stability result is placed in the appendix.
@@ 2.1. Spectral analysis
We describe the eigenvalues and eigenfunctions of the non-self-adjoint operators $L,L^{*}:D(L)=D(L^{*})=H^{2}(0,\pi)^{2}\cap H_{0}^{1}(0,\pi)^{2}\subset L^{2}(0,\pi)^{2}\rightarrow L^{2}(0,\pi)^{2}$ related to the following Sturm-Liouville problem
@@ 2.2. Controllability
It is worth mentioning that Lemma 2 was proved in [29, Theorem 1.2] for a forward system instead of a backward system. However, the above configuration is more appropriate for solving our inverse source problem (Problem 1). In addition, notice that the null controllability property with one scalar control for general complete matrices is not possible [29].
Concerning the system (1.2), the associated null controllability problem reads: given an initial datum $\Psi^{0}\in L^{2}(\Omega)^{n}$, we seek for a control function $U$ in $L^{2}(0,T;L^{2}(\mathcal{O})^{n})$ such that the corresponding solution $\Psi$ to
$\left\{\begin{array}[]{llll}-\partial_{t}\Psi-D^{*}\Delta\Psi=1_{\mathcal{O}}BU,&\mbox{in}&\Omega\times(0,T),\\ \Psi=0,&\mbox{on}&\partial\Omega\times(0,T),\\ \Psi(\,\cdot\,,T)=\Psi^{0},&\mbox{in}&\Omega,\end{array}\right$. (2.12)
satisfies the identity (2.10).
From [5, Remark 20], the system (2.12) can be controlled to zero with only one scalar control under the following assumptions (sufficient and necessary conditions):
  1. A1
    The diffusion matrix $D^{*}$ is diagonalizable with positive real eigenvalues, i.e., for $J=\operatorname{diag}(d_{i})_{n\times n}$ with $d_{1},d_{2},\dots,d_{n}>0$, one has $D^{*}=P^{-1}JP$, with $P\in\mathcal{M}_{n}(\mathbb{R})$, $\det P\neq 0$.
  2. A2
    $d_{i}\neq d_{j}$, for $i\neq j$,  $1\leq i,j\leq n$.
  3. A3
    $B=(b_{1},\dots,b_{n})^{t}\in\mathbb{R}^{n}$ and $b_{i}\neq 0$, for $i=1,\dots,n$.
The above can be summarized into the following Lemma.
Lemma 3.
For any $\Psi^{0}\in L^{2}(\Omega)^{n}$, system (2.12) is null controllable with one scalar control if and only if $D^{*}$ and $B$ satisfy A1, A2, and A3.
@@ 3. First main result: coupled systems with constant coefficients
In this section, we present explicit formulas for reconstructing the spectral coefficients of the vector source $F(x)$ in terms of a family of null controls $U^{(\tau)}$, which are linked to the systems (2.9) and (2.12), indexed by $\tau\in(0,T]$. In other words, we solve Problem 1. To achieve this, we begin by considering the following systems:
@@
and
$\left\{\begin{array}[]{llll}\partial_{t}Y-D\Delta Y=\sigma(t)F(x),&\mbox{in}&\Omega\times(0,T),\\ Y=0,&\mbox{on}&\partial\Omega\times(0,T),\\ Y(\,\cdot\,,0)=0,&\mbox{in}&\Omega,\end{array}\right$. (3.2)
where $Q\in L^{\infty}(\Omega)^{n^{2}}$ satisfies (2.7), (2.8), and $D$ satisfies  A1 and  A2.
@@
Let us observe that, thanks to the assumptions on the diffusion matrix $D$ and the potential matrix $Q$, for every source $\sigma(t)F(x)\in L^{2}(0,T;L^{2}(\Omega)^{n})$, system (3.1) (resp. system (3.2)) admits a unique weak solution $Y\in C([0,T];L^{2}(\Omega)^{n})\cap L^{2}(0,T;H^{1}_{0}(\Omega)^{n})$. Additionally, by considering $\sigma\in W^{1,\infty}(0,T)$ and following, for example, the procedure given in [33, Chapter 3, Section 6], solvability in the space
@@
also holds for systems (3.1) and (3.2).
@@
Before presenting the main results of this section, a common denominator to the systems (3.1) and (3.2) is the integral representation of their solutions, which are possible thanks to the linearity of those systems. More precisely, from the Duhamel principle, the solution $Y$ from either of those systems can be written by
@@
$\small{\left\{\begin{array}[]{llll}\partial_{t}W-\Delta W+QW=0,&\mbox{in}&\Omega\times(0,T),\\ W=0,&\mbox{on}&\partial\Omega\times(0,T),\\ W(\,\cdot\,,0)=\sigma(0)F(\cdot),&\mbox{in}&\Omega,\end{array}\right.\quad\mbox{or}\quad\left\{\begin{array}[]{llll}\partial_{t}W-D\Delta W=0,&\mbox{in}&\Omega\times(0,T),\\ W=0,&\mbox{on}&\partial\Omega\times(0,T),\\ W(\,\cdot\,,0)=\sigma(0)F(\cdot),&\mbox{in}&\Omega.\end{array}\right.}$ (3.4)
Furthermore, since $\partial_{t}Y(x,t)=\sigma(0)W(x,t)+\int\limits_{0}^{t}\partial_{t}\sigma(t-s)W(x,s)\,ds$, by evaluating at $t=T$ the main equations of (3.1) and (3.2), and using (3.3), we obtain the following identity:
@@
and
$\sigma(0)W(x,T)+\int\limits_{0}^{T}\partial_{t}\sigma(T-s)W(x,s)\,ds-D\Delta Y(x,T)=\sigma(T)F(x).$ (3.6)
As previously mentioned, our inverse source problems closely depend on the null controllability properties of the adjoint systems associated with (3.1) and (3.2), as well as on the spectral properties of the main operators. Consequently, the remainder of this section is dedicated to establishing explicit connections between the identities (3.5) and (3.6) and these essential theoretical components. Specifically, two key points must be addressed. First, the matrices $Q^{t},\,D^{*}\in\mathcal{M}_{n}(\mathbb{R})$, corresponding to the coupling matrices $Q,\,D\in\mathcal{M}_{n}(\mathbb{R})$ from systems (3.1) and (3.2), must satisfy the controllability conditions outlined in (2.8) and assumptions A1A2, and A3. Second, since the coupling matrices $Q$ and $D$ contain constant entries, we only need to consider $L^{2}$-eigenfunctions and eigenvalues of the Laplace operator with Dirichlet boundary conditions in $\Omega$ for the spectral assumptions. These eigenfunctions and eigenvalues will be denoted by $\{\varphi_{k}\}_{k\in\mathbb{N}}$ and $\{\lambda_{k}\}_{k\in\mathbb{N}}$, respectively. Finally, before stating our first result, we introduce some additional hypotheses.
@@
$a^{Q}_{{j,k}}(T):=\Biggl{(}1-\frac{\lambda_{k}}{\sigma(T)}\sum\limits_{i=1}^{n}m_{ij}(T)\Biggr{)}\neq 0,\quad\forall i,j=1,\dots,n,$ (3.7)
@@
$-(\Delta Y(T),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=-(Y(T),\Delta\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=\lambda_{k}(Y(T),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=\lambda_{k}\sum\limits_{i=1}^{n}(y_{i}(T),\varphi_{k})_{L^{2}(\Omega)}=\lambda_{k}\sum\limits_{i=1}^{n}y^{k}_{i}(T).$ (3.14)
@@
$+\frac{\lambda_{k}}{\sigma(T)}\sum\limits_{i=1}^{n}y^{k}_{i}(T)+\frac{1}{\sigma(T)}\int\limits_{0}^{T}\sigma(T-s)(QW(s),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}\,ds,\quad\sigma(T)\neq 0,$
@@
Now, we establish the necessary hypothesis for solving Problem 1 in the case where the coupling appears in the principal part, as is presented in the system (3.2).
H3.
$a^{D}_{{j,k}}(T):=\Biggl{(}1-\frac{\lambda_{k}}{\sigma(T)}\sum\limits_{\ell=1}^{n}\Biggl{(}\sum\limits_{i=1}^{n}d_{i\ell}\Biggr{)}m_{\ell j}(T)\Biggr{)}\neq 0,\quad\forall\,i,j=1,\dots,n,$ (3.27)
where $M=(m_{ij}(t))=\int\limits_{0}^{t}\tilde{\Phi}_{k}(t)\tilde{\Phi}_{k}^{-1}(s)\sigma(s)\,ds$ and $\tilde{\Phi}_{k}$ a fundamental matrix associated with the ordinary differential system: $Z^{\prime}+\lambda_{k}DZ=0$.
H4.
For every $k\in\mathbb{N}$ and $\Psi^{0}_{k}=(\varphi_{k},\dots,\varphi_{k})^{t}\in L^{2}(\Omega)^{n}$, consider Lemma 3 (null controllability property for (2.12) with datum $\Psi^{0}_{k},k\in\mathbb{N}$).
Our second source reconstruction result is given in the following theorem.
Theorem 2.
Let H3 and H4 be satisfied. Then, for every solution $Y\in W_{2}^{2,1}(\Omega\times(0,T))$ of (3.1), the source $F=(f_{1},\dots,f_{n})^{t}\in L^{2}(\Omega)^{n}$ satisfies the reconstruction identity
$\begin{array}[]{lll}\sum\limits_{j=1}^{n}a^{D}_{{j,k}}(T)(f_{j},\varphi_{k})_{L^{2}(\Omega)}&=-\displaystyle\frac{\sigma(0)}{\sigma(T)}(y_{n},\theta_{n})_{H^{1}(0,T;L^{2}(\mathcal{O}))}\\ &-\displaystyle\frac{1}{\sigma(T)}\int\limits_{0}^{T}\partial_{t}\sigma(T-s)(y_{n},\theta_{n})_{H^{1}(0,T;L^{2}(\mathcal{O}))}\,ds,\end{array}$ (3.28)
where $\theta_{n}$ is the unique solution to the Volterra equation (2.13) with right-hand side given by the control function obtained after solving the null controllability problem associated with the system (2.12).
Since the proof of Theorem 2 follows a guideline similar to the proof of Theorem 1, these details are presented in Appendix A.
@@ 4. Second main result: coupled systems with space dependent coefficients
$Q(x)=\begin{pmatrix}0&0\\ q(x)&0\end{pmatrix}\quad\mbox{and}\quad q(x)\geq C>0,\quad q\in L^{\infty}(0,\pi)\cap W^{1,\infty}(\tilde{\mathcal{O}}),\,\,\tilde{\mathcal{O}}\subset\mathcal{O}\subset(0,\pi).$
@@
$\left\{\begin{array}[]{llll}-\partial_{t}\Psi+L^{*}\Psi=(0,1_{\mathcal{O}}u_{2}),&\mbox{in }(0,\pi)\times(0,\tau),\\ \Psi(0,t)=\Psi(L,t)=0,&\mbox{in }(0,\tau),\\ \Psi(\,\cdot\,,\tau)=\Psi^{0}(\cdot),&\mbox{in }(0,\pi),\end{array}\right$. (4.3)
where $\Psi$ satisfies $\Psi(\,\cdot\,,0)=0\quad\mbox{in}\,\,(0,\pi)$ (i.e., system (4.3) satisfies the null controllability property, see [22, Theorem 1.1]).
@@
Let H5, H6, and H7 be satisfied. Let $\theta$ be a solution to (2.13) with right-hand side given by $1_{\mathcal{O}}u_{2}$. Then, for any solution $Y\in W_{2}^{2,1}((0,\pi)\times(0,T))$ of (4.1), the source $F=(f_{1},f_{2})^{t}\in L^{2}(0,\pi)^{2}$ satisfies
@@
$\mathcal{C}_{1}(\mathcal{B},\mathcal{B}^{*})=-\sigma(0)(y_{2},\theta)_{H^{1}(0,T;L^{2}(\mathcal{O}))},\quad\mathcal{C}_{2}(\mathcal{B},\mathcal{B}^{*})=-\displaystyle\int\limits_{0}^{T}\partial_{t}\sigma(T-s)(y_{2},\theta)_{H^{1}(0,T;L^{2}(\mathcal{O}))}\,ds,$
@@
Step 3. Controllability. As in the proof of Theorem 1, where global terms were replaced using two adjoint systems and their corresponding controllability results, we follow a similar strategy here. Specifically, the global inner-product terms $(\mathbf{w}(\,\cdot\,,\tau),\Phi_{1,k}^{*}+\Phi_{2,k}^{*})_{L^{2}(0,\pi)^{2}}$ for $\tau\in(0,T]$, are replaced by local terms in $L^{2}(0,\tau;L^{2}(\mathcal{O}))^{2}$, using the controllability assumption stated in Hypothesis H7.
@@
$a(T)(f_{1}^{\psi_{k}}+f_{2}^{\varphi_{k}}+f_{1}^{\varphi})+b(T)f_{1}^{\varphi}=-\sigma(0)(y_{2},\theta)_{H^{1}(0,T;L^{2}(\mathcal{O}))}-\displaystyle\int\limits_{0}^{T}\partial_{t}\sigma(T-s)(y_{2},\theta)_{H^{1}(0,T;L^{2}(\mathcal{O}))}\,ds,$

### v4

@@ Abstract.
This paper is devoted to the study of source reconstruction algorithms for coupled systems of heat equations, with either constant or spatially dependent coupling terms, where internal measurements are available from a reduced number of observed states. Two classes of systems are considered. The first comprises parabolic equations with constant zero-order coupling terms (through a matrix potential term). The second type considers parabolic equations coupled by a matrix potential that depends on spatial variables, which leads to the analysis of a non-self-adjoint operator. In all configurations, the source is assumed to be of separate variables, the temporal part is a known scalar function, and the spatial dependence is an unknown vector field. Several numerical examples using the finite element method in 1D and 2D are presented to show the reconstruction of space-dependent sources.
@@
Corresponding author - Cristhian Montoya: cdmontoyaz@eafit.edu.co
@@ 1.2. State of the art
    Inverse coefficient problems in coupled systems. Under the structure of systems (1.2) with $A=-\Delta$, the Bukhgeim-Klibanov method [29], which is based on Carleman estimates, has been widely employed to obtain results for inverse coefficient problems. For instance, the articles [15, 9] proved Lipschitz-type stability inequalities for $2\times 2$ reaction-diffusion systems using a single observation in a subdomain. These results were further extended to a nonlinear case in [14]. In a different context, the recent work [12] addressed the uniqueness and Lipschitz-type stability for cascade hyperbolic systems composed of $n$ equations, where internal measurements were available for all components of the solution except the last one. Another relevant work on inverse problems is [40], where the author presents self-contained derivations of Carleman estimates for parabolic equations and demonstrates their powerful applications in proving uniqueness and stability for inverse problems and control theory. Another recent work [17] considered two coupled Schrödinger equations (with $A=i\Delta$ in (1.2)), in which the authors investigated the logarithmic stability for determining two potentials from internal observations of a single component of the solution. In contrast to the previous results, [17] combined Carleman estimates along with the Fourier-Bros-Iagolnitzer transform in order to prove the logarithmic stability. The paper [13] tackled the problem of identification for two discontinuous coefficients in a one-dimensional coupled parabolic system, observing only one component. The authors proved theoretical results based on Carleman-type inequalities, while numerical results were obtained using the finite difference method for both temporal and spatial discretization, jointly with the interior-point method (meaning that the numerical part is treated as an optimization problem). To summarize, these articles show identifiability and stability results for hyperbolic and parabolic systems coupled through potential matrices (i.e., systems as (1.2)).
@@
The remaining of the paper is organized as follows: Section 3 provides an explicit source reconstruction formula for coupled parabolic systems with constant coefficient matrices (as a potential term into the system), thus solving Problem 1. Section 4 addresses Problem 2, presenting a reconstruction formula for a one-dimensional $2\times 2$ parabolic system coupled through a non-constant zero-order matrix potential. Here, the main challenge arises from the non-self-adjoint operator, necessitating careful use of eigenfunctions associated with non-self-adjoint Sturm–Liouville operators to achieve reconstruction based solely on interior measurements and a reduced number of states. Section 5 validates these theoretical findings through numerical experiments implemented with finite element methods and gradient descent algorithms. These experiments include reconstructions for Problem 2 in 1D, and reconstructions addressing Problem 1 in 2D. Finally, the Lipschitz-type stability result is placed in the appendix.
@@ 2.1. Spectral analysis
We describe the eigenvalues and eigenfunctions of the non-self-adjoint operators $L,L^{*}:H^{2}(0,\pi)^{2}\cap H_{0}^{1}(0,\pi)^{2}\subset L^{2}(0,\pi)^{2}\rightarrow L^{2}(0,\pi)^{2}$ related to the following Sturm-Liouville problem
@@ 2.2. Controllability
It is worth mentioning that Lemma 2 was proved in [26, Theorem 1.2] for a forward system instead of a backward system. However, the above configuration is more appropriate for solving our inverse source problem (Problem 1). In addition, notice that the null controllability property with one scalar control for general complete matrices is not possible [26].
@@ 3. First main result: coupled systems with constant coefficients
In this section, we present an explicit formula for reconstructing the spectral coefficients of the vector source $F(x)$ in terms of a family of null controls $U^{(\tau)}$, which are linked to the systems (2.9), indexed by $\tau\in(0,T]$. In other words, we solve Problem 1. To achieve this, we begin by considering the following system:
@@
where $Q\in L^{\infty}(\Omega)^{n^{2}}$ satisfies (2.7), (2.8).
@@
Let us observe that, thanks to the assumptions on the potential matrix $Q$, for every source $\sigma(t)F(x)\in L^{2}(0,T;L^{2}(\Omega)^{n})$, system (3.1) admits a unique weak solution $Y\in C([0,T];L^{2}(\Omega)^{n})\cap L^{2}(0,T;H^{1}_{0}(\Omega)^{n})$. Additionally, by considering $\sigma\in W^{1,\infty}(0,T)$ and following, for example, the procedure given in [30, Chapter 3, Section 6], solvability in the space
@@
also holds for system (3.1).
@@
Before presenting the main results of this section, we present integral representation to the solution for system (3.1), which are possible thanks to the linearity of this system. More precisely, from the Duhamel principle, the solution $Y$ from either of those systems can be written by
@@
$\small{\left\{\begin{array}[]{llll}\partial_{t}W-\Delta W+QW=0,&\mbox{in}&\Omega\times(0,T),\\ W=0,&\mbox{on}&\partial\Omega\times(0,T),\\ W(\,\cdot\,,0)=\sigma(0)F(\cdot),&\mbox{in}&\Omega,\end{array}\right.}$ (3.3)
Furthermore, since $\partial_{t}Y(x,t)=\sigma(0)W(x,t)+\int\limits_{0}^{t}\partial_{t}\sigma(t-s)W(x,s)\,ds$, by evaluating at $t=T$ the main equations of (3.1) and using (3.2), we obtain the following identity:
@@
As previously mentioned, our inverse source problems closely depend on the null controllability properties of the adjoint system associated with (3.1), as well as on the spectral properties of the main operator. Consequently, the remainder of this section is dedicated to establishing explicit connections between the identity (3.4) and this essential theoretical component. Specifically, two key points must be addressed. First, the matrix $Q^{t}\in\mathcal{M}_{n}(\mathbb{R})$, corresponding to the coupling matrices $Q,\in\mathcal{M}_{n}(\mathbb{R})$ from system (3.1), must satisfy the controllability conditions outlined in (2.8). Second, since the coupling matrix $Q$ contains constant entries, we only need to consider $L^{2}$-eigenfunctions and eigenvalues of the Laplace operator with Dirichlet boundary conditions in $\Omega$ for the spectral assumptions. These eigenfunctions and eigenvalues will be denoted by $\{\varphi_{k}\}_{k\in\mathbb{N}}$ and $\{\lambda_{k}\}_{k\in\mathbb{N}}$, respectively. Finally, before stating our first result, we introduce some additional hypotheses.
@@
$a^{Q}_{{j,k}}(T):=\Biggl(1-\frac{\lambda_{k}}{\sigma(T)}\sum\limits_{i=1}^{n}m_{ij}(T)\Biggr)\neq 0,\quad\forall\,j=1,\dots,n,$ (3.5)
@@
$-(\Delta Y(T),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=\lambda_{k}(Y(T),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}=\lambda_{k}\sum\limits_{i=1}^{n}(y_{i}(T),\varphi_{k})_{L^{2}(\Omega)}=\lambda_{k}\sum\limits_{i=1}^{n}y^{k}_{i}(T).$ (3.12)
@@
$+\frac{\lambda_{k}}{\sigma(T)}\sum\limits_{i=1}^{n}y^{k}_{i}(T)+\frac{1}{\sigma(T)}\int\limits_{0}^{T}\sigma(T-s)(QW(s),\Psi^{0}_{k})_{L^{2}(\Omega)^{n}}\,ds,\qquad\sigma(T)\neq 0,$
@@ 4. Second main result: coupled systems with space dependent coefficients
$Q(x)=\begin{pmatrix}0&0\\ q(x)&0\end{pmatrix}\quad\mbox{and}\,\,q(x)\geq C>0,\quad q\in L^{\infty}(0,\pi)\cap W^{1,\infty}(\tilde{\mathcal{O}}),\,\,\tilde{\mathcal{O}}\subset\mathcal{O}\subset(0,\pi).$
@@
$\left\{\begin{array}[]{llll}-\partial_{t}\Psi+L^{*}\Psi=(0,1_{\mathcal{O}}u_{2}),&\mbox{in }(0,\pi)\times(0,\tau),\\ \Psi(0,t)=\Psi(\pi,t)=0,&\mbox{in }(0,\tau),\\ \Psi(\,\cdot\,,\tau)=\Psi^{0}(\cdot),&\mbox{in }(0,\pi),\end{array}\right$. (4.3)
where $\Psi$ satisfies $\Psi(\,\cdot\,,0)=0\quad\mbox{in}\,\,(0,\pi)$ (i.e., system (4.3) satisfies the null controllability property, see [19, Theorem 1.1]).
@@
Let H3, H4, and H5 be satisfied. Let $\theta$ be a solution to (2.12) with right-hand side given by $1_{\mathcal{O}}u_{2}$. Then, for any solution $Y\in W_{2}^{2,1}((0,\pi)\times(0,T))$ of (4.1), the source $F=(f_{1},f_{2})^{t}\in L^{2}(0,\pi)^{2}$ satisfies
@@
$\mathcal{C}_{1}(\mathcal{B},\mathcal{B}^{*})=-\sigma(0)(y_{2},\theta)_{H^{1}(0,T;L^{2}(\mathcal{O}))},$
$\mathcal{C}_{2}(\mathcal{B},\mathcal{B}^{*})=-\displaystyle\int\limits_{0}^{T}\partial_{t}\sigma(T-s)(y_{2},\theta)_{H^{1}(0,T;L^{2}(\mathcal{O}))}\,ds,$
@@
Step 3. Controllability. As in the proof of Theorem 1, where global terms were replaced using two adjoint systems and their corresponding controllability results, we follow a similar strategy here. Specifically, the global terms $(\mathbf{w}(\,\cdot\,,\tau),\Phi_{1,k}^{*}+\Phi_{2,k}^{*})_{L^{2}(0,\pi)^{2}}$ for $\tau\in(0,T]$, are replaced by local terms in $L^{2}(0,\tau;L^{2}(\mathcal{O}))^{2}$, using the controllability assumption stated in Hypothesis H5.
@@
$a(T)(f_{1}^{\psi_{k}}+f_{2}^{\varphi_{k}}+f_{1}^{\varphi})+b(T)f_{1}^{\varphi}$ $=-\sigma(0)(y_{2},\theta)_{H^{1}(0,T;L^{2}(\mathcal{O}))}$
$\quad-\int\limits_{0}^{T}\partial_{t}\sigma(T-s)(y_{2},\theta)_{H^{1}(0,T;L^{2}(\mathcal{O}))}\,ds,$